OpenStudy (anonymous):

Using the formula C= 5/9(F - 32), find the Fahrenheit temperature when the Celsius temperature is 45º.

OpenStudy (zpupster):

remember this one it is easier f=1.8C + 32 F=1.8(45) + 32 F= 113 we can do it your way also 45= 5/9(F-32) mult both sides by 9/5 45(9/5) = F-32 81=F-32 add 32 both sides 113 = F

OpenStudy (anonymous):

Thank you so much c:

OpenStudy (zpupster):

you are welcome!